OpenStudy (needshelpdfdf):

The solution set for -p2 − 11p = 0 is

OpenStudy (eliesaab):

\[ -p^2-11p=0\\ -p(p+11)=0 \]

OpenStudy (needshelpdfdf):

hmmm

OpenStudy (needshelpdfdf):

idk how to do this like how do I work it out

OpenStudy (eliesaab):

The product of two factors is zero when at least one of them is zero

OpenStudy (eliesaab):

\[ -p=0 \implies p=0\\ p+11=0 \implies p=-11 \]